OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 10%. What percent of the original volume is the volume of the new prism?

OpenStudy (anonymous):

Do you know the formula for volume of Rectangular Prism??

OpenStudy (anonymous):

\[Volume = Length \times Width \times Height\]

OpenStudy (anonymous):

find the length and width of rectangular base after increment find volume before and after increament and calculate percentage

OpenStudy (anonymous):

volume of rectangular prism is 1/3(area of base times height)

OpenStudy (anonymous):

So now each dimension is increased by 10%. Say length be l l is increased by 10% So now it has becomes: l + \(\frac{10}{100}l\) = \(\frac{11}{10}l\) Similarly other two lengths will be the same: \[Volume = \frac{11 \times 11 \times 11}{1000}LWH\]

OpenStudy (anonymous):

This is the new volume..